Bob is a vector with magnitude 240 and direction 200 degrees. Mai is a vector of length 160 @ 310 degrees. Jose is a vector of 210 @ 170 degrees. Calculate the following vectors and report their magnitude and angle (measured counterclockwise from the positive x-axis):

(a) Bob + Mai

(b) Mai – Bob

(c) Jose + Bob – Mai

Comment: "If possible. I'd like to know the how and why on some steps, please. Thanks"
Answer #1

Given : Bob : 240@200° , Mai: 160@310° , Jose: 210@170°

Solution:

(a) Bob + Mai

Components of vector Bob : Bobx = 240cos(200)= -225.53

Boby= 240sin(200) = -82.08

Components of vector Mai: Maix = 160cos(310)= 102.85

Maiy = 160sin(310)= -122.57

Vector addition in x direction :

Bobx+Maix = -225.53 +102.85 = -122.68

Vector addition in y direction:

Boby +Maiy = -82.08 -122.57 = -204.65

Magnitude of vector addition |Bob +Mai| is given by:

\sqrt[(-122.68)2+(-204.65)2] = 238.60

Direction of vector addition is given by:

\theta = tan-1(-204.65/-122.68) = 59.06°

Hence

Bob + Mai = 238.60 @ 59.06°

(b) Mai - Bob

Vector subtraction in x direction:

Maix - Bobx = 102.85 -(-225.53) = 328.38

Vector subtraction in y direction:

Maiy - Boby = -122.57 -(-82.08) = -40.49

Magnitude of vector subtraction (|Mai - Bob|) is given by:

\sqrt[(328.38)2+(-40.49)2] = 330.87

Direction of vector subtraction is given by:

\theta = tan-1(-40.49/328.38) = -7.02° (clockwise)

\theta = 360-7.02= 352.97° (anticlockwise)

Hence

Mai - Bob = 330.87@352.97°

(c) Jose + Bob - Mai

Vector components of Jose : Josex = 210cos(170)= -206.81

Josey = 210sin(170)= 36.47

X direction:

Josex+Bobx-Maix = -206.81 -225.53 -102.85 = -121.57

y direction:

Josey+Boby-Maiy= 36.47-82.08-(-122.57) = 76.96

Magnitude | Jose+Bob-Mai| is given by:

\sqrt[(-122.57)2+(76.96)2] = 144.73

Its direction is given by:

\theta = tan-1(76.96/-121.57) = -32.34° (clockwise)

\theta = 360°-32.34° = 327.66° (anticlockwise)

Hence

Jose + Bob - Mai = 144.73@ 327.66°